OpenStudy (anonymous):

I have never seen this before how do i go about doing it? solve for x in terms of pi sin(x + pi/4)+sin(x - pi/4)=-1

OpenStudy (anonymous):

do i need to use the sin identities? sin(A)cos(b)+sin(B)cos(A) and sin(A)cos(b)-sin(B)cos(A)?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

thanks for that ill post the answer when i get it

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

see sin(x + pi/4) +sin(x - pi/4)=(1/(v2))(sinx+cosx) +(1/(v2))(sinx-cosx)=V2 sinx hence V2sinx=-1 or sinx=-1/sqrt(2) = or sinx =-sin(pi/4) or sinx=sn(pi+pi/4) =sin(5pi/4) hence x=5pi/4

OpenStudy (anonymous):

Why can't you let him do it? @matricked

OpenStudy (anonymous):

lol dude when helping people on this site @matricked you should rather let them do it themselves rather than providing the answer. There is no benefit in doing that.

OpenStudy (lgbasallote):

you still have a chance to do it yourself @Razzputin ..matricked didn't use the trig formulas

OpenStudy (anonymous):

yup i know im just asking him to refrain from doing what he does, when someone does that my eyes just drift over.

OpenStudy (lgbasallote):

spoils the fun doesn't it

OpenStudy (anonymous):

so i get to \[2\sin \frac{ \sqrt{2} }{ 2 }+2\cos \frac{ \sqrt{2} }{ 2 }=-1\] am i right so far?

OpenStudy (anonymous):

\[sinx+cosx=\frac{ -1 }{ \sqrt{2} }\]

OpenStudy (anonymous):

i think i did something wrong?

OpenStudy (anonymous):

Yes..recheck.

OpenStudy (anonymous):

\[sinx * \cos \frac{ \pi }{ 4 }+\sin\frac{ \pi }{ 4 } * cosx +sinx * \cos -\frac{ \pi }{ 4 }-\sin-\frac{ \pi }{ 4 } * cosx=-1\] \[sinx *\frac{ \sqrt2 }{ 2 }+\frac{ \sqrt2 }{ 2 } * cosx +sinx * \frac{ \sqrt2 }{ 2 }--\frac{ \sqrt2 }{ 2 } * cosx=-1\] \[\frac{ \sqrt2 }{ 2 }sinx +\frac{ \sqrt2 }{ 2 }cosx +\frac{ \sqrt2 }{ 2 }sinx+\frac{ \sqrt2 }{ 2 }cosx=-1\] \[2*\frac{ \sqrt2 }{ 2 }sinx +2*\frac{ \sqrt2 }{ 2 }cosx=-1\] \[ \sqrt2 sinx +\sqrt2 cosx=-1\] \[ sinx +cosx=\frac{ -1 }{ \sqrt2 }\] where did i go wrong?

OpenStudy (anonymous):

\[\sin(x+\pi/4) = sinx.\cos \pi/4 + cosx.\sin \pi/4\] \[\sin(x-\pi/4) = sinx.\cos \pi/4 - cosx.\sin \pi/4\]

OpenStudy (anonymous):

You made a mistake in your first step. did you notice it now?

OpenStudy (anonymous):

ahhh i see so the sign does not cary into the equation rather determines which equation?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

If you carry the sign into the equation, you have to use only the first equation...

OpenStudy (anonymous):

so i need to have \[-1 = 2sinx \frac{\sqrt2}{2}\]

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

so x = -45??

OpenStudy (anonymous):

\[-\frac{1}{4}\pi\]

OpenStudy (anonymous):

The general solution is \[(4n+3)\pi/2 \pm \pi/4\]

OpenStudy (anonymous):

the exact answer depends on the interval given

OpenStudy (anonymous):

its not right. i get 0 instead of -1 if i sub in\[ -\frac{1}{4}\pi\] how do you get to the answer you got?

OpenStudy (anonymous):

No If you substitute -45 in your question, you get -1...

OpenStudy (anonymous):

but -45 is the same as /[-frac{1}{4}pi/]

OpenStudy (anonymous):

the answer is \[\sin x = -1/\sqrt{2}\]

OpenStudy (anonymous):

Hey wait...what are you saying? Let us be clear..What did you get and what is wrong?

OpenStudy (anonymous):

Yes -45 is same as -pi/4

OpenStudy (anonymous):

ok i got \[sinx = \frac{ -1 }{ \sqrt2 }\] which equals x = -45 in degrees and \[x= \frac{-1}{4}\pi\] in radians. When i substitute those two answers in i dont get -1

OpenStudy (anonymous):

sin(-45+45) + sin(-45-45) = sin0 + sin(-90) = 0 + (-1) = -1

OpenStudy (anonymous):

ok now type that in but instead use the values x = -pi/4 and pi/4

OpenStudy (anonymous):

i get - 0.02741213359

OpenStudy (anonymous):

thats if i use my calculator set to degrees once i set it to radians it gives me -1...

OpenStudy (anonymous):

thank you for your help @AbhimanyuPudi

OpenStudy (anonymous):

welcome:)

OpenStudy (anonymous):

cool, i learnt two things in one here... never mix degrees and radians on a calculator it isnt clever enough to determine that and also the trick you showed me with the two equations